Prove that any two open intervals of the real.line are homeomorphic


1
Expert's answer
2021-07-29T11:47:03-0400

Solution:

Proof:

Let's show that a bounded open interval (a, b) is homeomorphic to (0,1).

Consider the map f:(0,1) "\rightarrow" (a, b) defined by f(t)=a+(b-a) t. It is continuous and bijective. It's inverse is continuous as well.

Now let's show that (0,1) is homeomorphic to (a, "\infty" ). Consider g:(0,1) "\rightarrow(a, \infty)" defined by

"g(t)=a+\tan (\pi t / 2)"

This is clearly bijective and continuous. Its inverse is continuous as well.

It's obvious that (-"\infty", a) is homeomorphic to (-a, "\infty").


Now we have that (-1,1) is homeomorphic to R using "h(t)=\tan (\pi t / 2)"

The property of being homeomorphic is transitive.

Hence, any two open intervals of the real line are homeomorphic.